# Tight Algorithm for Connected Odd Cycle Transversal Parameterized by Clique-width

### 12 Feb 2024 v1

@@ 5.1 Action sequences
Each node $y\in V(\mathcal{T}_{x})$ induces an action-subsequence $\tau_{y}$ of $\tau$ at $y$ (denoted by $\tau_{y}$) given by the restriction of $\tau$ to $\mathcal{V}^{CJ}_{y}$. Each action sequence $\tau$ at a node $x$ generates a pattern $p^{\tau}_{x}\in\mathscr{P}_{C}$ and a coloring $c^{\tau}_{x}\in\overline{C}^{[k]}$ (defined next). We say that $\tau$ generates the pair $(p^{\tau}_{x},c^{\tau}_{x})$. For $y\in\mathcal{V}_{x}$, we denote $p^{\tau_{y}}_{y}$ by $p^{\tau}_{y}$, and $c^{\tau_{y}}_{y}$ by $c^{\tau}_{y}$. We call an action-sequence $\tau$ at a node $x$ valid, if and only if for each join node $y\in\mathcal{V}_{x}$, it holds that $c^{\tau}_{y}(i)$ and $c^{\tau}_{y}(j)$ are consistent.
@@
  • Create node $\mu_{x}=i(v)$: For $X\in\{\boldsymbol{\emptyset},\boldsymbol{B},\boldsymbol{W}\}$ we define $c^{i}_{X}:[k]\rightarrow\overline{C}$ as
@@
Given two families $S,R\subseteq\mathscr{P}$, We say that $R$ represents $S$, if for each $q\in\mathscr{P}$ the following holds: there exists a pattern $p\in S$ such that $p\sim q$ if and only if there exists a pattern $p^{\prime}\in R$ such that $p^{\prime}\sim q$.
@@
The operations defined in Definition 3.2 as well as the operation $\square_{i,j}$ defined in Definition 3.3 for all $i,j\in[k]$ with $i\neq j$ preserve representation.
@@
Given a labeled graph $G=(V,E)$ and a set of vertices $S$, we denote with $\operatorname{pat}_{G}(S)$ the pattern defined as follows: let $C_{0}=\emptyset$ if $v_{0}\notin S$, or $C_{0}$ be the connected component of $G[S]$ containing $v_{0}$. Let $C_{1},\dots C_{r}$ be the connected components of $G[S]$ not containing $v_{0}$.
@@
In the following, we show that one can represent any partial solution by action-sequences. In order to achieve this we need the following definition:
@@
Each node $y\in V(\mathcal{T}_{x})$ induces a solution-subsequence $\pi_{y}$ of $\pi$ at $y$ given by the restriction of $\pi$ to $\mathcal{V}^{C}_{y}$. Each solution sequence $\pi$ at a node $x$ generates a pattern $o^{\pi}_{x}\in\mathscr{P}$ and a coloring $c^{\pi}_{x}\in\overline{C}^{[k]}$ (defined next). We say that $\pi$ generates the pair $(o^{\pi}_{x},c^{\pi}_{x})$. For $y\in V(\mathcal{T}_{x})$, we denote $o^{\pi_{y}}_{y}$ by $o^{\pi}_{y}$, and $c^{\pi_{y}}_{y}$ by $c^{\pi}_{y}$. We call a solution-sequence $\tau$ at a node $x$ valid, if and only if for each join node $y\in V(\mathcal{T}_{x})$, it holds that that $c^{\tau}_{y}(i)$ and $c^{\tau}_{y}(j)$ are consistent.
We define $(p^{\tau}_{y},c^{\tau}_{y})$ recursively as follows:
  • Create node $\mu_{x}=i(v)$: For $X\in\{\boldsymbol{\emptyset},\boldsymbol{B},\boldsymbol{W}\}$ we define $c_{X}:[k]\rightarrow\overline{C}$ as
@@
Let $x\in\mathcal{V}$, and $S\subseteq V_{x}$. Let $g:V_{x}\setminus S\rightarrow\{\boldsymbol{B},\boldsymbol{W}\}$ be some mapping. Let $\pi=\pi^{S,g}_{x}:\mathcal{V}^{C}_{x}\rightarrow\{0,3,4\}$ be the solution-sequence at $x$ defined as follows: for $y\in\mathcal{V}^{C}_{x}$, let $v\in V_{x}$ be the vertex created at $y$, i.e. $\mu_{y}=i(v)$ for some $i\in[k]$. We define
@@
Let $x\in\mathcal{V}$. For $S\subseteq V_{x}$ and some mapping $g:V_{x}\setminus S\rightarrow\{\boldsymbol{B},\boldsymbol{W}\}$, let $\pi=\pi^{S,g}_{x}$. Then it holds that $\operatorname{pat}_{x}(S)=o^{\pi}_{x}$, $\operatorname{col}_{x}(g)=c^{\pi}_{x}$ and that $g$ is a proper 2-coloring of $V_{x}\setminus S$ in $G_{x}$ if and only if $\pi$ is a valid solution-sequence at $x$.
@@
We prove by induction over $\mathcal{T}_{x}$ that for $y\in\mathcal{V}_{x}$ the following holds
  1. 1.
    $\operatorname{pat}_{y}(S)=o^{\pi}_{y}$.
  2. 2.
    $\operatorname{col}_{y}(g)=c^{\pi}_{y}$.
  3. 3.
    $g_{y}$ is a proper coloring of $G_{y}$ if and only if $c_{y}$ is a valid solution-subsequence.
We distinguish different types of nodes $y$. We begin with the base case:
  • Create node $\mu_{y}=i(v)$: If $v\in S$, it holds that $\pi(y)=0$. Hence, it holds for $p=[0i]$ if $v=v_{0}$, or $p=[i]$ otherwise, that $\operatorname{pat}_{y}(S)=o^{\pi}_{y}=p$, and that $\operatorname{col}_{y}(g)(j)=c^{\pi}_{y}(j)=\boldsymbol{\emptyset}$ for $j\in[k]$.
    Assuming that $v\notin S$, it holds that $\pi(y)\neq 0$. Hence, it holds that $\operatorname{pat}_{y}(S)=o^{\pi}_{y}=[0]$. It holds that $\pi(y)=3$ if $g(v)=\boldsymbol{B}$ and $\pi(y)=4$ otherwise. In both cases, it holds that $\operatorname{col}_{y}(g)(i)=c^{\pi}_{y}(i)=g(v)$, and $\operatorname{col}_{y}(g)(j)=c^{\pi}_{y}(j)=\boldsymbol{\emptyset}$ for all $j\in[k]\setminus\{i\}$.
@@
  • Relabel node $\mu_{y}=\rho_{i\rightarrow j}(\mu_{y^{\prime}})$: Let $p=\operatorname{pat}_{y^{\prime}}(S)$, and $c=\operatorname{col}_{y^{\prime}}(g)$. It holds by induction hypothesis that $o^{\pi}_{y^{\prime}}=p$ and $c=c^{\pi}_{y^{\prime}}$. The graph $G_{y}$ results from $G_{y^{\prime}}$ by changing the label of each $i$-vertex in $G_{y^{\prime}}$ into $j$. Hence, it holds clearly that
@@
  • Join node $\mu_{y}=\eta_{i,j}(\mu_{y^{\prime}})$: Let $p=\operatorname{pat}_{y^{\prime}}(S)$, and $c=\operatorname{col}_{y^{\prime}}(g)$. It holds by induction hypothesis that $o^{\pi}_{y^{\prime}}=p$ and $c=c^{\pi}_{y^{\prime}}$. The graph $G_{y}$ results from $G_{y^{\prime}}$ by adding all edges between vertices of label $i$ and vertices of label $j$, turning them into one connected components if vertices of both labels exist in $G_{y}$ or keeping $G_{y^{\prime}}$ as it is otherwise. Hence, it holds clearly that
@@
    • $o^{\pi}_{y_{1}}=\operatorname{pat}_{y_{1}}(S)=p_{1}$
    • $o^{\pi}_{y_{2}}=\operatorname{pat}_{y_{2}}(S)=p_{2}$
    • $c^{\pi}_{y_{1}}=\operatorname{col}_{y_{1}}(g)=c_{1}$.
    • $c^{\pi}_{y_{2}}=\operatorname{col}_{y_{2}}(g)=c_{2}$.
    Let $p=p_{1}\oplus p_{2}$, and $c=c_{1}\oplus_{C}c_{2}$. It follows from Definition 5.8, that $o^{\pi}_{y}=p$ and $c^{\pi}_{y}=c$.
    The graph $G_{y}$ results from the disjoint union of $G_{y_{1}}$ and $G_{y_{2}}$. Hence, the connected components of $G_{y}$ is the union of their connected components. Since $v_{0}$ can belong to at most one of $V_{y_{1}}$ and $V_{y_{2}}$, it holds as well that the zero-set of at least one of $p_{1}$ and $p_{2}$ is a singleton. Hence, it holds clearly that $\operatorname{pat}_{y}(S)=p_{1}\oplus p_{2}=p$.
@@
Let $x\in\mathcal{V}$, $q\in\mathscr{P}$ and $c\in\mathscr{C}$. There exists a set of vertices $S\subseteq V_{x}$ of size ${b}$ and a mapping $g:V_{x}\setminus S\rightarrow\{\boldsymbol{B},\boldsymbol{W}\}$ of $G_{x}$ such that $\operatorname{pat}_{x}(S)\sim q$, and $c=\operatorname{col}_{x}(g)$, if and only if there exists a complete pattern $p^{\prime}\in\mathscr{P}_{C}$ with $p^{\prime}\sim q$, a weight ${w}$ and an action-sequence $\tau$ of cost ${b}$ and weight ${w}$ at $x$ generating the pair $(p^{\prime},c)$. Moreover, $g$ is a proper 2-coloring of $G_{x}\setminus S$ if and only if $\tau$ is valid.
@@
Let $\pi:\mathcal{V}^{C}_{x}\rightarrow\{0,3,4\}$ be an arbitrary solution sequence at $x$, and let $(p,c)$ be the pair generated by $\pi$. Let $\Pi=\Pi_{x}(\pi)$ be the family of all action-sequences at $x$, where for a create node $y\in\mathcal{V}^{C}_{x}$, it holds that $\tau(x)=\pi(x)$ if $\tau(x)\in\{3,4\}$, or $\tau(x)\in\{1,2\}$ if $\pi(x)=0$. We define the family of patterns $P=P_{x}(\pi)=\{p^{\tau}_{x}\colon\tau\in\Pi\}$. We show by induction over $\mathcal{T}$ that the family $P_{x}(\pi)$ represents $p$, and that $c^{\tau}_{x}=c$ for all $\tau\in\Pi$. Moreover, for each $\tau\in\Pi$ it holds that $\tau$ is valid if and only if $\pi$ is.
Assuming this is the case. Let $S\subseteq V_{x}$, and $g:V_{x}\setminus S\rightarrow\{\boldsymbol{B},\boldsymbol{W}\}$. For $p=\operatorname{pat}_{x}(S)$, $c=\operatorname{col}_{x}(g)$ and $\pi=\pi^{S,g}_{x}$, it holds by Lemma 5.10 that $o^{\pi}_{x}=p$, $c^{\pi}_{x}=c$ and that $g$ is a proper 2-coloring of $V_{x}\setminus S$ in $G_{x}$ if and only if $\pi$ is a valid solution-sequence at $x$. By the claim above, it holds that $p\sim q$ if and only if there exists a pattern $p^{\prime}\in P$ with $p^{\prime}\sim q$. Hence, it holds that $p\sim q$ if and only if there exists an action-sequence $\tau\in\Pi(\pi)$ at $x$ generating the pair $(p^{\prime},c)$ with $p^{\prime}\sim q$, and $\tau$ is a valid action-sequence at $x$ if and only if $g$ is a proper two-coloring of $G_{x}\setminus S$.
Now we prove the claim by induction over $\mathcal{T}_{v}$. We assume that $\pi$ generates $(p,c)$ at $x$, and $\Pi=\Pi_{x}(\pi)$, and $P=P_{x}(\pi)$.
  • Create node $\mu_{x}=i(v)$: If $\pi(x)=0$, it holds for $p=o^{\pi}_{x}$, that $P=R_{p}$. It follows from Lemma 5.4 that $P$ represents $p$. It holds for $\Pi=\{\tau_{1},\tau_{2}\}$ that all sequences $\pi,\tau_{1},\tau_{2}$ generate the coloring $c$ with $c(j)=\boldsymbol{\emptyset}$ for all $j\in[k]$. On the other hand, if $\pi(x)\in\{3,4\}$, it holds that $o^{\pi}_{x}=[0]$, and $P=\{[0]\}$, so $P$ trivially represents. In this case $\Pi$ is a singleton $\tau$ with $\tau(x)=\pi(x)$. Hence, it holds that $c^{(}_{\tau})=c^{(}_{\pi})=c$, where $c(j)=\boldsymbol{\emptyset}$ for all $j\in[k]\setminus\{i\}$, and $c(i)=\boldsymbol{B}$ if $\pi(x)=3$ or $c(i)=\boldsymbol{W}$ otherwise. Finally, it holds that all mentioned sequences are valid ones.
@@
    It holds that $o^{\pi}_{x}=\square_{i,j}o^{\pi_{x^{\prime}}}_{x^{\prime}}$. Let $\hat{P}=\square_{i,j}P^{\prime}$. By induction hypothesis, it holds that $P^{\prime}$ represents $p^{\prime}$, and hence, by Lemma 5.6 it holds that $\hat{P}$ represents $p$. From the definition of $p^{\tau}_{x}$, it holds that
    $P=\bigcup\limits_{q\in\hat{P}}R_{q}.$
    By Lemma 5.4 it holds for each $q\in\hat{P}$ that there exists some set $Q\subseteq P$ that represents $q$. Hence, $P$ represents $\hat{P}$, and by transitivity of representation, $P$ represents $p$ as well.
@@
    It follows that $o^{\pi}_{x}=o^{\pi_{x_{1}}}_{x_{1}}\oplus o^{\pi_{x_{2}}}_{x_{2}}$ and from Definition 5.1 $P=P_{1}\oplus P_{2}$. By induction hypothesis, it holds that $P_{1}$ represents $p_{1}$, and $P_{2}$ represents $p_{2}$. Hence, by Lemma 5.6 it holds that $P$ represents $p$. Also, by induction hypothesis, it holds for $i\in\{1,2\}$, and for each $\tau_{i}\in P_{i}$ that $\tau_{i}$ generates the coloring $c_{i}$ at $x_{i}$ and that $\tau_{i}$ is valid if and only if $\pi_{i}$ is.
@@
Let $x\in\mathcal{V}$, and $S\subseteq V_{x}$ be a non-empty set of vertices. Then it holds that $\operatorname{pat}_{x}(S)\sim[0]$ if and only if $G_{x}[S]$ is connected, and $v_{0}\in S$.
@@
Since $S$ is not empty, clearly $G_{x}[S]$ contains a connected component that does not contain $v_{0}$ if $v_{0}\notin S$ or $G_{x}[S]$ is not connected, which is the case if and only if $p$ contains a set differing from the zero-set, and hence $p$ is not a singleton. The claim follows from the fact that $p\sqcup[0]=p$ for all patterns $p$. ∎

### 26 Feb 2024 v2

@@ 5.1 Action-sequences
Each node $y\in V(\mathcal{T}_{x})$ induces an action-subsequence $\tau_{y}$ of $\tau$ at $y$ (denoted by $\tau_{y}$) given by the restriction of $\tau$ to $\mathcal{V}^{CJ}_{y}$. Each action-sequence $\tau$ at a node $x$ generates a pattern $p^{\tau}_{x}\in\mathscr{P}_{C}$ and a coloring $c^{\tau}_{x}\in\overline{C}^{[k]}$ (defined next). We say that $\tau$ generates the pair $(p^{\tau}_{x},c^{\tau}_{x})$. For $y\in\mathcal{V}_{x}$, we denote $p^{\tau_{y}}_{y}$ by $p^{\tau}_{y}$, and $c^{\tau_{y}}_{y}$ by $c^{\tau}_{y}$. We call an action-sequence $\tau$ at a node $x$ valid, if and only if for each join node $y\in\mathcal{V}_{x}$, where $\mu_{y}=\eta_{i,j}(\mu_{y^{\prime}})$, it holds that that $c^{\tau}_{y^{\prime}}(i)$ and $c^{\tau}_{y^{\prime}}(j)$ are consistent.
@@
  • Introduce node $\mu_{x}=i(v)$: For $X\in\{\boldsymbol{\emptyset},\boldsymbol{B},\boldsymbol{W}\}$ we define $c^{i}_{X}:[k]\rightarrow\overline{C}$ as
@@
Given two families $S,R\subseteq\mathscr{P}$, we say that $R$ represents $S$, if for each $q\in\mathscr{P}$ the following holds: there exists a pattern $p\in S$ such that $p\sim q$ if and only if there exists a pattern $p^{\prime}\in R$ such that $p^{\prime}\sim q$. We say that the family $S$ represents a pattern $p\in\mathscr{P}$, if $S$ represents $\{p\}$.
Observation 5.3.
Given four families $S_{1},S_{2},R_{1},R_{2}\subseteq\mathscr{P}$, where $R_{1}$ represents $S_{1}$, and $R_{2}$ represents $S_{2}$, it holds that $R_{1}\cup R_{2}$ represents $S_{1}\cup S_{2}$.
@@
The operations join, relabel and union, as defined in Definition 3.2, as well as the operation $\square_{i,j}$ defined in Definition 3.3 for all $i,j\in[k]$ with $i\neq j$ preserve representation.
@@
Given a labeled graph $G=(V,E)$ and a set of vertices $S$, we denote with $\operatorname{pat}_{G}(S)$ the pattern defined as follows: let $C_{0}=\emptyset$ if $v_{0}\notin S$, or let $C_{0}$ be the connected component of $G[S]$ containing $v_{0}$ otherwise. Let $C_{1},\dots C_{r}$ be the connected components of $G[S]$ not containing $v_{0}$. Then we define
@@
In the following, we show that one can represent any partial solution by action-sequences. In order to achieve this we introduce the notion of solution-sequences. Similar to an action-sequence, a solution-sequence is defined by operations over $\mathcal{T}$. However, a solution-sequence only indicates actions, that specify which vertices are included in a partial solution, and hence, it corresponds to a whole solution, instead of a representing pattern thereof.
@@
Each node $y\in V(\mathcal{T}_{x})$ induces a solution-subsequence $\pi_{y}$ of $\pi$ at $y$ given by the restriction of $\pi$ to $\mathcal{V}^{C}_{y}$. Each solution-sequence $\pi$ at a node $x$ generates a pattern $o^{\pi}_{x}\in\mathscr{P}$ and a coloring $c^{\pi}_{x}\in\overline{C}^{[k]}$ (defined next). We say that $\pi$ generates the pair $(o^{\pi}_{x},c^{\pi}_{x})$. For $y\in V(\mathcal{T}_{x})$, we denote $o^{\pi_{y}}_{y}$ by $o^{\pi}_{y}$, and $c^{\pi_{y}}_{y}$ by $c^{\pi}_{y}$. We call a solution-sequence $\tau$ at a node $x$ valid, if and only if for each join node $y\in\mathcal{V}_{x}$, where $\mu_{y}=\eta_{i,j}(\mu_{y^{\prime}})$, it holds that $c^{\tau}_{y^{\prime}}(i)$ and $c^{\tau}_{y^{\prime}}(j)$ are consistent.
We define $(o^{\tau}_{y},c^{\tau}_{y})$ recursively as follows:
  • Introduce node $\mu_{x}=i(v)$: For $X\in\{\boldsymbol{\emptyset},\boldsymbol{B},\boldsymbol{W}\}$ we define $c_{X}:[k]\rightarrow\overline{C}$ as
@@
Let $x\in\mathcal{V}$, and $S\subseteq V_{x}$. Let $g:V_{x}\setminus S\rightarrow\{\boldsymbol{B},\boldsymbol{W}\}$ be some mapping. Let $\pi=\pi^{S,g}_{x}:\mathcal{V}^{C}_{x}\rightarrow\{0,3,4\}$ be the solution-sequence at $x$ defined as follows: for $y\in\mathcal{V}^{C}_{x}$, let $v\in V_{x}$ be the vertex introduced at $y$, i.e. $\mu_{y}=i(v)$ for some $i\in[k]$. We define
@@
Let $x\in\mathcal{V}$. For $S\subseteq V_{x}$ and some mapping $g:V_{x}\setminus S\rightarrow\{\boldsymbol{B},\boldsymbol{W}\}$, let $\pi=\pi^{S,g}_{x}$. Then all the following holds:
  1. 1.
    $b(\pi)=|S|$,
  2. 2.
    $\operatorname{pat}_{x}(S)=o^{\pi}_{x}$,
  3. 3.
    $\operatorname{col}_{x}(g)=c^{\pi}_{x}$,
  4. 4.
    and $g$ is a proper 2-coloring of $G_{x}[V_{x}\setminus S]$ if and only if $\pi$ is valid.
@@
First, we prove Item 1. It holds that
$b(\pi)=|\{y\in\mathcal{V}^{C}_{x}\colon\pi(y)=0\}|=|v\in S|,$
where the first equality holds from Definition 5.9, and the second from Definition 5.10. Now we prove the other points for all $y\in\mathcal{V}_{x}$ by induction over $\mathcal{T}_{x}$. We distinguish the different types of nodes $y$.
  • Introduce node $\mu_{y}=i(v)$ (base case): If $v\in S$, then it holds that $\pi(y)=0$. Hence, it holds for $p=[0i]$ if $v=v_{0}$, or $p=[i]$ otherwise, that $\operatorname{pat}_{y}(S)=o^{\pi}_{y}=p$. It also holds that $\operatorname{col}_{y}(g)(j)=c^{\pi}_{y}(j)=\boldsymbol{\emptyset}$ for $j\in[k]$.
    Now we assume that $v\notin S$. It follows that $\pi(y)\neq 0$. Hence, it holds that $\operatorname{pat}_{y}(S)=o^{\pi}_{y}=[0]$. It holds that $\pi(y)=3$ if $g(v)=\boldsymbol{B}$ and $\pi(y)=4$ otherwise. In both cases, it holds that $\operatorname{col}_{y}(g)(i)=c^{\pi}_{y}(i)=g(v)$, and $\operatorname{col}_{y}(g)(j)=c^{\pi}_{y}(j)=\boldsymbol{\emptyset}$ for all $j\in[k]\setminus\{i\}$.
@@
  • Relabel node $\mu_{y}=\rho_{i\rightarrow j}(\mu_{y^{\prime}})$: Let $p=\operatorname{pat}_{y^{\prime}}(S)$, and $c=\operatorname{col}_{y^{\prime}}(g)$. It holds by induction hypothesis that $o^{\pi}_{y^{\prime}}=p$ and $c=c^{\pi}_{y^{\prime}}$. The graph $G_{y}$ results from $G_{y^{\prime}}$ by changing the label of each vertex labeled $i$ in $G_{y^{\prime}}$ to the label $j$. Hence, it holds clearly that
@@
  • Join node $\mu_{y}=\eta_{i,j}(\mu_{y^{\prime}})$: Let $p=\operatorname{pat}_{y^{\prime}}(S)$, and $c=\operatorname{col}_{y^{\prime}}(g)$. It holds by induction hypothesis that $o^{\pi}_{y^{\prime}}=p$ and $c=c^{\pi}_{y^{\prime}}$. The graph $G_{y}$ results from $G_{y^{\prime}}$ by adding all edges between vertices of label $i$ and vertices of label $j$, turning them into one connected component if vertices of both labels exist in $G_{y}$, or by keeping $G_{y^{\prime}}$ as it is otherwise. Hence, it holds clearly that
@@
    • $o^{\pi}_{y_{1}}=\operatorname{pat}_{y_{1}}(S)=p_{1}$,
    • $o^{\pi}_{y_{2}}=\operatorname{pat}_{y_{2}}(S)=p_{2}$,
    • $c^{\pi}_{y_{1}}=\operatorname{col}_{y_{1}}(g)=c_{1}$,
    • and $c^{\pi}_{y_{2}}=\operatorname{col}_{y_{2}}(g)=c_{2}$.
    Let $p=p_{1}\oplus p_{2}$, and $c=c_{1}\oplus_{C}c_{2}$. It follows from Definition 5.9, that $o^{\pi}_{y}=p$ and $c^{\pi}_{y}=c$. The graph $G_{y}$ results from the disjoint union of $G_{y_{1}}$ and $G_{y_{2}}$. Hence, the connected components of $G_{y}$ consist of the union of the connected components of both graphs $G_{y_{1}}$ and $G_{y_{2}}$. Since $v_{0}$ can belong to at most one of $V_{y_{1}}$ and $V_{y_{2}}$, it holds as well that the zero-set of at least one of $p_{1}$ and $p_{2}$ is a singleton. If follows that $\operatorname{pat}_{y}(S)=p_{1}\oplus p_{2}=p$.
@@
Let $x\in\mathcal{V}$, $q\in\mathscr{P}$ and $c\in\mathscr{C}$. There exists a set of vertices $S\subseteq V_{x}$ of size ${b}$ and a mapping $g:V_{x}\setminus S\rightarrow\{\boldsymbol{B},\boldsymbol{W}\}$, such that $\operatorname{pat}_{x}(S)\sim q$, and $c=\operatorname{col}_{x}(g)$, if and only if there exists a complete pattern $p^{\prime}\in\mathscr{P}_{C}$ with $p^{\prime}\sim q$, a weight ${w}$ and an action-sequence $\tau$ of cost ${b}$ and weight ${w}$ at $x$ generating the pair $(p^{\prime},c)$. Moreover, $g$ is a proper 2-coloring of $G_{x}\setminus S$ if and only if $\tau$ is valid.
@@
Let $\pi:\mathcal{V}^{C}_{x}\rightarrow\{0,3,4\}$ be an arbitrary solution-sequence at $x$, and let $(p,c)$ be the pair generated by $\pi$. Let $\Pi=\Pi_{x}(\pi)$ be the family of all action-sequences $\tau$ at $x$, such that for each introduce node $y\in\mathcal{V}^{C}_{x}$, it holds that $\tau(x)=\pi(x)$ if $\tau(x)\in\{3,4\}$, or $\tau(x)\in\{1,2\}$ if $\pi(x)=0$. We define the family of patterns $P=P_{x}(\pi)=\{p^{\tau}_{x}\colon\tau\in\Pi\}$. By induction over $\mathcal{T}$, we will prove that all following points hold:
  1. 1.
    The family $P_{x}(\pi)$ represents $p$.
  2. 2.
    It holds for all $\tau\in\Pi$ that $c^{\tau}_{x}=c$.
  3. 3.
    It holds for all $\tau\in\Pi$ that $\tau$ is valid if and only if $\pi$ is.
  4. 4.
    It holds for all $\tau\in\Pi$ that $b(\tau)=b(\pi)$.
Before we prove this claim, we show that the lemma follows from it. So we assume first that the claim holds. Let $S\subseteq V_{x}$ be a set of vertices, and $g:V_{x}\setminus S\rightarrow\{\boldsymbol{B},\boldsymbol{W}\}$ be some mapping. Let $p=\operatorname{pat}_{x}(S)$, $c=\operatorname{col}_{x}(g)$ and $\pi=\pi^{S,g}_{x}$. It holds by Lemma 5.11 that $o^{\pi}_{x}=p$, $c^{\pi}_{x}=c$, $b(\pi)=|S|$, and that $g$ is a proper 2-coloring of $V_{x}\setminus S$ in $G_{x}$ if and only if $\pi$ is a valid solution-sequence at $x$.
Hence, it holds for $q\in\mathscr{P}$ by the claim above (Item 1), that $p\sim q$ if and only if there exists a pattern $p^{\prime}\in P$ with $p^{\prime}\sim q$. Hence, it holds that $p\sim q$ if and only if there exists an action-sequence $\tau\in\Pi_{x}(\pi)$ of cost $|S|$ at $x$ generating the pair $(p^{\prime},c)$ for some pattern $p^{\prime}$ with $p^{\prime}\sim q$. Finally, it holds by Item 3 that $\tau$ is a valid action-sequence at $x$ if and only if $g$ is a proper two-coloring of $G_{x}\setminus S$.
We first prove Item 4. It holds that
$b(\tau)$ =|$=\big{|}\big{\{}y\in\mathcal{V}^{C}_{x}\colon\tau(y)\in\{1,2\}\big{\}}\big{|}$
=|$=|\{y\in\mathcal{V}^{C}_{x}\colon\pi(y)=0\}|$ $=b(\pi),$
where the first equality follows from Definition 5.1, the second from the definition of $\Pi_{x}(\pi)$, and the third from Definition 5.9. Now we prove the other points by induction over $\mathcal{T}$. Let $\pi$ be some solution-sequence at $x$ of cost ${b}$, and let $(p,c)$ be the pair generated by $\pi$. Let $\Pi=\Pi_{x}(\pi)$, and $P=P_{x}(\pi)$.
  • Introduce node $\mu_{x}=i(v)$: If $\pi(x)=0$, it holds for $p=o^{\pi}_{x}$, that $P=R_{p}$ (Definition 5.4). It follows from Lemma 5.5 that $P$ represents $p$. It holds for $\Pi=\{\tau_{1},\tau_{2}\}$ that all sequences $\pi,\tau_{1},\tau_{2}$ generate the coloring $c$ with $c(j)=\boldsymbol{\emptyset}$ for all $j\in[k]$. On the other hand, if $\pi(x)\in\{3,4\}$, it holds that $o^{\pi}_{x}=[0]$, and $P=\{[0]\}$, so $P$ trivially represents $p$. In this case $\Pi$ is a singleton $\tau$ with $\tau(x)=\pi(x)$. Hence, it holds that $c^{\tau}_{x}=c^{\pi}_{x}=c$, where $c(j)=\boldsymbol{\emptyset}$ for all $j\in[k]\setminus\{i\}$, and $c(i)=\boldsymbol{B}$ if $\pi(x)=3$ or $c(i)=\boldsymbol{W}$ otherwise. Finally, it holds that all mentioned sequences are valid ones.
@@
    It holds that $o^{\pi}_{x}=\square_{i,j}o^{\pi_{x^{\prime}}}_{x^{\prime}}$. Let $\hat{P}=\square_{i,j}P^{\prime}$. By induction hypothesis, it holds that $P^{\prime}$ represents $p^{\prime}$, and hence, by Lemma 5.7 it holds that $\hat{P}$ represents $p$. Since it holds from Definition 5.4 for $q\in\hat{P}$, that $R_{q}=\{\operatorname{Ac}(q,i)\colon i\in[4]\}$, it holds from the definition of $p^{\tau}_{x}$ (Definition 5.1) that
    $P=\bigcup\limits_{\ell\in[4]}\operatorname{Ac}(\hat{P},\ell)=\bigcup\limits_{q\in\hat{P}}R_{q}.$
    Finally, it holds by Lemma 5.5 and 5.3 that $P$ represents $\hat{P}$, and hence, by transitivity of representation, that $P$ represents $p$ as well.
@@
    It follows that $o^{\pi}_{x}=o^{\pi_{x_{1}}}_{x_{1}}\oplus o^{\pi_{x_{2}}}_{x_{2}}$ and from Definition 5.1, that $P=P_{1}\oplus P_{2}$. By induction hypothesis, it holds that $P_{1}$ represents $p_{1}$, and $P_{2}$ represents $p_{2}$. Hence, by Lemma 5.7 it holds that $P$ represents $p$. Also, by the induction hypothesis, it holds for each $i\in\{1,2\}$, and for each $\tau_{i}\in P_{i}$ that $\tau_{i}$ generates the coloring $c_{i}$ at $x_{i}$ and that $\tau_{i}$ is valid if and only if $\pi_{i}$ is.
@@
Let $x\in\mathcal{V}$, and $S\subseteq V_{x}$ be a non-empty set of vertices. Then it holds that $p=\operatorname{pat}_{x}(S)\sim[0]$ if and only if $G_{x}[S]$ is connected, and $v_{0}\in S$.
@@
It holds that $p\sim[0]$ if and only if $p$ consists of the zero-set only. However, since $S$ is not empty, it holds that $G_{x}[S]$ contains at least one connected component. Hence, $p$ consists of the zero-set only, if and only if $v_{0}\in S$ and $G_{x}[S]$ is connected. ∎